Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-page---20130303

Observers: Nic Walton, Nadia Blagorodnova, Matthew Nicholl

Sunset: UT 23:21 / Twilight 00:35 - 09:14 / Sunrise 10:28      LST 6-16hrs

Classification      Followup    Stds:     Tech probs

All times in this report approximate UT   (Chilean time currently UT - 3hrs)

EFOSC: All OBs in the 188.D-3003(Q)/VM/EFOSC2 directory and 
SOFI: All OBs in the 188.D-3003(R)/VM/SOFI directory on the observers OB Execution machine (wg5dhs)

On wg5dhs - OBs held in /home/visitor/p2pp-impex/nblago 

Phot std finders at http://www.eso.org/sci/observing/tools/standards/Landolt.html

Afternoon Calibrations

UT: 1850 - arc-spec   (note OB in 188.D-3003(P) for this one)
UT: 1920 - flat-spec
UT: 1945 - flat-JHK - all SOFI calibration obtained

Twilight Calibrations

None

SCIENCE OBSERVATIONS (SOFI)

Start at UT: 2345 - lots of cirrus cloud in the north, some elsewhere. 

UT: 2345 - IR spec std - GD71 - in B and G gratings. (OB execution time 10m)
60s/120s exp - std looks very faint, especially in the red. will re do later

UT: 2355 - image analysis on mirror

UT: 0010 - LSQ12hxg (> 19mag)  JHK (1h45m)

UT: 0100 - sky looks clear - cloud gone.

UT: 0145 - Telluric standard for PSNJ09 - Hip045705  (V = 7.96, I = 7.26) ( 5m )

UT: 0155 - PSNJ09491471-474456 BG (1h40m x 1.35)
Marshal requests just BG for this. 

UT: 0430 cloud getting heavier 

UT: 0440 - Telluric standard for PSNJ09 - Hip045705  (V = 7.96, I = 7.26) ( 5m )

UT: 0445 - Spec Std  SOFI LTT 3218 (10 * 1.35 = m)

UT: 0505 - PSNJ09491471-474456 JHK (35m * 1.35) 
(no guiding as guide star not visible due to cloud)

SCIENCE OBSERVATIONS (EFOSC2)

Change to EFOSC2

UT: 0600 - arcs

UT: 0610 - PSNJ13540068-0755438 - gr13  (400s exp - due to cloud)
(cloud - but seeing good at ~0.7)
SNID: classified as a Ia at z=0.009 - epoch pre max. 

UT: 0645 - CSS130303105205-133424 - gr13 (1500s exp - due to cloud) 
SNID: classified as SN 1a 91T - 5-7d premax - z=0.07
Note: this target taken from the INBOX as no suitable high priority candidates in the classify queue. 

UT: 0720 - PSNJ09491471-474456 - gr13 (OB exc time 25m)

UT: 0740 - Flux Std EFOSC EG274

UT: 0800 - SN2013K grism13 (35min) 
Finder at SN2013K_FC.png - note that the coords of the SN refer to the centre of the host galaxy - not the SN

UT: 0845 - had to stop as too cloudy - no guide stars

UT: 0855 - arcs - PESSTO_calib_only_arcs

UT: 0910 - end of night - didn't look like cloud would clear. 

seeing stayed good ~0.7 all night. 











Comments